# The odd independence number of graphs, I: Foundations and classical classes

### v1

@@ Abstract
We prove that $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds for all claw-free graphs $G$, and present many results, using various techniques, concerning the odd independence number of cycles, paths, Moore graphs, Kneser graphs, the complete subdivision $S(K_{n})$ of $K_{n}$, the half graphs $H_{n,n}$, and $K_{p}\,\Box\,K_{q}$. Further, we consider the odd independence number of the hypercube $Q_{d}$ and also of the complements of triangle-free graphs. Many open problems for future research are stated.
@@ 1 Introduction
An odd independent set $S$ in a graph $G=(V,E)$ is an independent set of vertices such that, for every vertex $v\in V\,\diagdown\,S$, either $N(v)\cap S=\varnothing$ or $|N(v)\cap S|\equiv 1$ (mod 2), where $N(v)$ stands for the open neighbourhood of $v$. The largest cardinality of odd independent sets in $G$, denoted $\alpha_{\mathrm{od}}(G)$, is called the odd independence number of $G$. This new parameter is a natural companion to the recently introduced strong odd chromatic number [3, 7, 9, 10, 11]. A proper vertex coloring of a graph $G$ is a strong odd coloring if, for every vertex $v\in V(G)$, each color used in the neighborhood of $v$ appears an odd number of times in $N(v)$. The minimum number of colors in a strong odd coloring of $G$ is denoted by $\chi_{\mathrm{so}}(G)$. This paper is aimed at introducing and starting a systematic study of the odd independence number.
In Section 2 we present basic inequalities relating $\alpha_{\mathrm{od}}(G)$, $\chi_{\mathrm{so}}(G)$ and other graphs parameters like $|G|$ (the order of $G$) and $\Delta(G)$ (the maximum degree of $G$), similar to such classical inequalities involving $\alpha(G)$ and $\chi(G)$. We also supply upper bounds for $\alpha_{\mathrm{od}}(G)$ for $d$-regular graphs when $d$ is even, or when $G$ is a $K_{1,r}$-free graph, or when any two adjacent vertices in $G$ have at least $\lambda$ vertices in common. We also mention that $\alpha_{\mathrm{od}}(G)=\alpha(G)$ holds for every regular bipartite graph $G$ with odd degree, and we further prove $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ for claw-free graphs $G$, complementing the result [3] that $\chi_{\mathrm{so}}(G)=\chi(G^{2})$ holds for claw-free graphs. Then we present various lower bounds concerning $\alpha_{\mathrm{od}}(G)$ using known results on $\chi_{\mathrm{so}}(G)$ and the inequality $\alpha_{\mathrm{od}}(G)\geq|G|/\chi_{\mathrm{so}}(G)$, which serves as a benchmark for further results.
@@ 6 Complements of triangle-free graphs
Such graphs exists, the smallest one is the self-complementary $G=P_{4}$ with $\alpha_{\mathrm{od}}(G)=\alpha(G)=2$, cf. Problem 33 $(i)$. Also the graph $G=C_{6}=v_{1}v_{2}\dots v_{6}$ supplemented with the edges $(v_{1},v_{4})$ and $(v_{2},v_{5})$ is triangle-free of diameter 3. Its complement consists of two triangles connected by an edge, hence having diameter 3 and $\alpha_{\mathrm{od}}(\overline{G})=\alpha(\overline{G})=2$ (but $\alpha(G)=3$).
More generally, consider $H=K_{n,m}-e$ (one edge omitted), with $n\geq m\geq 2$. This $H$ is triangle-free and has $\mathrm{diam}(H)=3$, forced by the vertices of the removed edge. Its complement $G=\overline{H}$ is $K_{n}\cup K_{m}$ supplemented with one edge connecting the two complete parts, hence having diameter 3. Further, $\alpha_{\mathrm{od}}(G)=\alpha(G)=2$ (cf. again Problem 33).
@@
In this case $\alpha_{\mathrm{od}}(\overline{G})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$ holds (cf. Problem 35).
@@ 8 Concluding remarks and open problems
    Characterize the graphs $G$ for which $\alpha_{\mathrm{od}}(G)=\alpha(G)$ holds.
@@
    Characterize the graphs for which $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds.
@@
Problem 35.
Characterize graphs $G$ with $\chi_{\mathrm{so}}(G)=|G|$, or equivalently with $\alpha_{\mathrm{od}}(G)=1$; cf. Proposition 1.4.1 (lower bound) and 1.4.2 (upper bound).
@@
Problem 37.
Determine the computational complexity of decision problems associated with Problems 33, 35, and 36.

### v2

@@ Abstract
We prove that $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds for all claw-free graphs $G$, and apply this result to prove that determining $\alpha_{\mathrm{od}}(G)$ is in general NP-hard (and also when restricted to line graphs). We also present many results, using various techniques, concerning the odd independence number of cycles, paths, Moore graphs, Kneser graphs, the complete subdivision $S(K_{n})$ of $K_{n}$, the half graphs $H_{n,n}$, and $K_{p}\,\Box\,K_{q}$. Further, we consider the odd independence number of the hypercube $Q_{d}$ and also of the complements of triangle-free graphs. Many open problems for future research are stated. Further related results can be found in part II of this work [The odd independence number of graphs, II: Finite and infinite grids and chessboard graphs, arXiv: 2510.01897].
@@ 1 Introduction
An odd independent set $S$ in a graph $G=(V,E)$ is an independent set of vertices such that, for every vertex $v\in V\,\diagdown\,S$, either $N(v)\cap S=\varnothing$ or $|N(v)\cap S|\equiv 1$ (mod 2), where $N(v)$ stands for the open neighbourhood of $v$. The largest cardinality of odd independent sets in $G$, denoted $\alpha_{\mathrm{od}}(G)$, is called the odd independence number of $G$. This new parameter is a natural companion to the recently introduced strong odd chromatic number [3, 7, 11, 13, 14]. A proper vertex coloring of a graph $G$ is a strong odd coloring if, for every vertex $v\in V(G)$, each color used in the neighborhood of $v$ appears an odd number of times in $N(v)$. The minimum number of colors in a strong odd coloring of $G$ is denoted by $\chi_{\mathrm{so}}(G)$. This paper is aimed at introducing and starting a systematic study of the odd independence number; further results on $\alpha_{\mathrm{od}}$ are given in the follow-up manuscript [4].
In Section 2 we present basic inequalities relating $\alpha_{\mathrm{od}}(G)$, $\chi_{\mathrm{so}}(G)$ and other graphs parameters like $|G|$ (the order of $G$) and $\Delta(G)$ (the maximum degree of $G$), similar to such classical inequalities involving $\alpha(G)$ and $\chi(G)$. We also supply upper bounds for $\alpha_{\mathrm{od}}(G)$ for $d$-regular graphs when $d$ is even, or when $G$ is a $K_{1,r}$-free graph, or when any two adjacent vertices in $G$ have at least $\lambda$ vertices in common. We also mention that $\alpha_{\mathrm{od}}(G)=\alpha(G)$ holds for every regular bipartite graph $G$ with odd degree. Moreover, concerning the class of claw-free graphs $G$, complementing the result $\chi_{\mathrm{so}}(G)=\chi(G^{2})$ from [3] we prove that also $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds for them, and apply this identity to prove that determining $\alpha_{\mathrm{od}}(G)$ is NP-hard, even when restricted to many types of line graphs. Then we present various lower bounds concerning $\alpha_{\mathrm{od}}(G)$ using known results on $\chi_{\mathrm{so}}(G)$ and the inequality $\alpha_{\mathrm{od}}(G)\geq|G|/\chi_{\mathrm{so}}(G)$, which serves as a benchmark for further results.
@@ 2.1 Forbidden pair lemma
An important consequence is the following complexity result.
Theorem 11.
The problem of determining $\alpha_{\mathrm{od}}(G)$ is NP-hard. It also remains NP-hard when the input graph is restricted (among others) to the line graphs $L(H)$ of graphs from any of the following graph classes: planar graphs of maximum degree $4$, bipartite graphs of maximum degree $3$, $C_{4}$-free bipartite graphs, $r$-regular graphs with $r\geq 5$, line graph graphs, chair-free graphs, or Hamiltonian graphs.
Proof.
We apply reduction from the problem Maximum Induced Matching (MIM), which asks, for a generic input graph $H$, to determine the largest number of edges no two of which are adjacent or incident in $H$. Induced matching in $H$ precisely means a set of vertices in the line graph $L(H)$ which have pairwise distances at least 3, hence an independent set in the graph $(L(H))^{2}$. Every line graph is claw-free, thus denoting $G:=L(H)$ we have $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ by Theorem 10. The number of steps in this reduction from $H$ to $G$ and $G^{2}$ is polynomial in terms of input size. As a consequence, if MIM is intractable on a certain class of graphs $H$, then so is the determination of $\alpha_{\mathrm{od}}$ on the line graphs $L(H)$ of those graphs. NP-hardness of MIM on the graph classes listed in our theorem follows from the results proved in the references [9], [10], [12]. $\Box$
@@ 6 Complements of triangle-free graphs
Such graphs exists, the smallest one is the self-complementary $G=P_{4}$ with $\alpha_{\mathrm{od}}(G)=\alpha(G)=2$, cf. Problem 34 $(i)$. Also the graph $G=C_{6}=v_{1}v_{2}\dots v_{6}$ supplemented with the edges $(v_{1},v_{4})$ and $(v_{2},v_{5})$ is triangle-free of diameter 3. Its complement consists of two triangles connected by an edge, hence having diameter 3 and $\alpha_{\mathrm{od}}(\overline{G})=\alpha(\overline{G})=2$ (but $\alpha(G)=3$).
More generally, consider $H=K_{n,m}-e$ (one edge omitted), with $n\geq m\geq 2$. This $H$ is triangle-free and has $\mathrm{diam}(H)=3$, forced by the vertices of the removed edge. Its complement $G=\overline{H}$ is $K_{n}\cup K_{m}$ supplemented with one edge connecting the two complete parts, hence having diameter 3. Further, $\alpha_{\mathrm{od}}(G)=\alpha(G)=2$ (cf. again Problem 34).
@@
In this case $\alpha_{\mathrm{od}}(\overline{G})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$ holds (cf. Problem 36).
@@ 8 Concluding remarks and open problems
    Characterize the graphs $G$ for which $\alpha_{\mathrm{od}}(G)=\alpha(G)$ holds; or determine the complexity of deciding if $\alpha_{\mathrm{od}}(G)=\alpha(G)$.
@@
    Characterize the graphs for which $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds; or determine the complexity of deciding if $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$.
@@
Problem 36.
Characterize graphs $G$ with $\chi_{\mathrm{so}}(G)=|G|$, equivalently with $\alpha_{\mathrm{od}}(G)=1$; or determine the complexity of the corresponding decision problem.